A cube shaped jewelry box has a surface area of 300 square inches,What are the dimensions of the jewelry box?
Answer by Alan3354(69443)

A cube has 6 faces of equal area
Area = 6s^2 (s = side length)
6s^2 = 300
s^2 = 50
s = sqrt(50) = 5sqrt(2)
s =~ 7.0711 inches